7. Carbon absorbs energy at a wavelength of 150.0 nm. The total amount of energy emitted by a carbon sample is 1.98 x 105 J. Calculate the number of carbon atoms present in the sample, assuming that each atom emits one photon.


1
Expert's answer
2022-04-19T03:11:58-0400

The first thing we need to do is to convert the wavelength to meters. We should remember that 1 meter is 1×109 nm so:

"\lambda" = 150 nm / 1×109 nm = 1.5×10-7 m

The formula for Energy in this case is:

E = hc/"\lambda"

We already know that c is the speed of light, which is 3×108 m/s and h is the planck constant which is 6.63×10-34 J s

Replacing data in the above formula:

E = 6.63×10-34 × 3×108 / 1.5×10-7 = 1.33×10-18 J per photon.

Now finally to calculate the carbon atoms present in the sample, let's calculate the number of photons:

N of photons = 1.98×105 / 1.33×10-18

N of photons = 1.49×1023 photons

The exercise states that one atom emits on photon, so we can assume that this final result would be the number of carbon atoms present in the sample.